### v1

@@ 4 Application to the QROT problem
To apply our proposed ripALM, we first reformulate (4.1) as follows:
$\min_{X\in\mathbb{R}^{m\times n}}~{}\Big{\{}f_{\texttt{q}}(X):=\frac{\lambda}{2}\left\|{X}\right\|_{F}^{2}+\left\langle{C},{X}\right\rangle+\delta_{\mathbb{R}^{m\times n}_{+}}(X)\;\mid\;X\bm{1}_{n}=\bm{\alpha},~{}X^{\top}\bm{1}_{m}=\bm{\beta}\Big{\}}.$ (4.2)
Clearly, it falls into the form of problem (1.1) (upon vectorization of $X$). Further, one can show that the dual problem of (4.2) is given by (modulo a minus sign)
@@
and its associated augmented Lagrangian function is given by (using the similar deduction as Section 2)
@@
$X^{k+1}=\texttt{prox}_{\sigma_{k}f_{\texttt{q}}}\left(X^{k}+\sigma_{k}\left(\bm{u}^{k+1}\bm{1}_{n}^{\top}+\bm{1}_{m}(\bm{v}^{k+1})^{\top}\right)\right),$ (4.5)
@@
To truly implement our ripALM for solving problem (4.2), it is essential to efficiently solve the subproblem (4.4) to find $(\bm{u}^{k+1},\bm{v}^{k+1})$ satisfying the error criterion (2.5). In the following, we shall describe how to apply a semismooth Newton (Ssn) method to achieve this goal. For simplicity, we drop the index $k$ and explicitly rewrite the subproblem (4.4) as follows:
$\min\limits_{\bm{u}\in\mathbb{R}^{m},\bm{v}\in\mathbb{R}^{n}}\left\{\begin{aligned} &\Psi(\bm{u},\bm{v}):=-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\frac{1}{2\sigma}\left\|\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right\|_{F}^{2}-\frac{1}{2\sigma}\|\bar{X}\|_{F}^{2}\\ &\quad-\mathtt{M}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)+\frac{\tau}{2\sigma}\left\|\bm{u}-\bar{\bm{u}}\right\|^{2}+\frac{\tau}{2\sigma}\left\|\bm{v}-\bar{\bm{v}}\right\|^{2}\end{aligned}\right\},$ (4.6)
@@
$\nabla\Psi(\bm{u},\bm{v})=\left[\begin{matrix}\left(\texttt{prox}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)\right)\bm{1}_{n}-\bm{\alpha}+\tau\sigma^{-1}\left(\bm{u}-\bar{\bm{u}}\right)\vspace{1mm}\\ \left(\texttt{prox}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)\right)^{\top}\bm{1}_{m}-\bm{\beta}+\tau\sigma^{-1}\left(\bm{v}-\bar{\bm{v}}\right)\end{matrix}\right].$
Note that the proximal mapping $\texttt{prox}_{\sigma f_{\texttt{q}}}(\cdot)$ can be easily computed as follows:
$\texttt{prox}_{\sigma f_{\texttt{q}}}\left(Z\right)=\frac{1}{1+\lambda\sigma}\Pi_{\mathbb{R}_{+}^{m\times n}}\left(Z-\sigma{C}\right),\quad\forall\,Z\in\mathbb{R}^{m\times n}.$
@@
In view of the nice property of $\nabla\Psi$, we are able to follow [19, 21, 22, 51] to apply a globally convergent and locally superlinearly convergent Ssn method to solve (4.7). Specifically, let $\mathcal{W}(\bm{u},\bm{v}):=\bar{X}+\sigma\left({\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-C}\right)$ and define the multifunction ^$\widehat{\partial}(\nabla\Psi):\mathbb{R}^{m}\times\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{(m+n)\times(m+n)}$ as follows:
^$\widehat{\partial}(\nabla\Psi)(\bm{u},\bm{v}):=\left\{H\in\mathbb{R}^{(m+n)\times(m+n)}~{}\left\lvert~{}\begin{aligned} H:=&\;\frac{\sigma}{1+\lambda\sigma}\,B\mathrm{Diag}(\texttt{vec}(\Omega))B^{\top}+\frac{\tau}{\sigma}I_{m+n},\\[3.0pt] \forall&\,\Omega\in\partial{\Pi_{\mathbb{R}_{+}^{m\times n}}}(\mathcal{W}(\bm{u},\bm{v}))\end{aligned}\right.\right\},$
where $B:=\begin{bmatrix}\bm{1}_{n}^{\top}\otimes I_{m}\\ I_{n}\otimes\bm{1}_{m}^{\top}\end{bmatrix}\in\mathbb{R}^{(m+n)\times mn}$ with $\otimes$” denoting the Kronecker product, $\texttt{vec}(\Omega)$ denotes the vectorization of $\Omega$ with $[\texttt{vec}(\Omega)]_{i+(j-1)m}=\Omega_{ij}$ for any $1\leq i\leq m$ and $1\leq j\leq n$, $\mathrm{Diag}(\bm{z})$ denotes the diagonal matrix whose $i$th diagonal element is given by $z_{i}$, and $\partial{\Pi_{\mathbb{R}_{+}^{m\times n}}}(Z)$ denotes the generalized Jacobian of the Lipschitz continuous mapping $\Pi_{\mathbb{R}_{+}^{m\times n}}$ at $Z$, which is defined by
@@
Then, using similar arguments as in the proof of [51, Proposition 4], one can show that $\nabla\Psi(\cdot)$ is strongly semi-smooth with respect to ^$\widehat{\partial}(\nabla\Psi)(\cdot)$, and thus the Ssn method is applicable. More importantly, for any $(\bm{u},\bm{v})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, all elements of ^$\widehat{\partial}(\nabla\Psi)(\bm{u},\bm{v})$ are positive definite. This ensures the direct applicability of the Ssn method without the need for a specific regularity condition, such as the primal constraint nondegeneracy condition [55], highlighting the advantage of incorporating a proximal term. The detailed description of the Ssn method for solving equation (4.7) is presented in Algorithm 2. We refer readers to [19, Theorem 3.6] for its detailed convergence results.
@@ 5 Numerical experiments
In this section, we conduct numerical experiments to evaluate the performance and validate the efficiency of our ripALM in Algorithm 1 for solving the QROT problem (4.1). Specifically, we will conduct experiments from the following two aspects:
  • In Section 5.1, we compare our ripALM with Snipal [21] and ciPALM [51], which represent the latest absolute-type and relative-type inexact pALMs, respectively. These comparisons aim to illustrate how different error criteria with different tolerance parameters influence the practical numerical performance of the inexact pALM.
  • In Section 5.2, we compare our ripALM with Gurobi and a dual alternating direction method of multipliers (dADMM, see Appendix B for more details) for solving large-scale QROT problems. Gurobi represents one of the state-of-the-art commercial solvers, while the dADMM is a widely used first-order method for solving large-scale constrained convex optimization problems; see, for example, [5, 14].
@@
Termination conditions. Let $\mathcal{Z}(\bm{u},\bm{v},X):=C+\lambda X-\bm{u}\bm{1}_{n}^{\top}-\bm{1}_{m}\bm{v}^{\top}$. The Karush-Kuhn-Tucker (KKT) system for problem (4.1) and its dual (4.3) is given by
@@
$\Delta_{\rm kkt}(\bm{u},\bm{v},X):=\max\left\{\Delta_{p}(X),\,\Delta_{d}(\bm{u},\bm{v},X),\,\Delta_{c}(\bm{u},\bm{v},X)\right\},$
@@
[…]
@@
$\Delta_{\rm gap}(\bm{u},\,\bm{v},\,X):=\cfrac{\left|\mathtt{pobj}(X)-\mathtt{dobj}(\bm{u},\bm{v})\right|}{1+\left|\mathtt{pobj}(X)\right|+\left|\mathtt{dobj}(\bm{u},\bm{v})\right|},$
where $\mathtt{pobj}(X):=f_{\texttt{q}}(X)$ and $\mathtt{dobj}(\bm{u},\bm{v}):=-f_{\texttt{q}}^{*}\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)+\bm{\alpha}^{\top}\bm{u}+\bm{\beta}^{\top}\bm{v}$. Using these relative residuals, we will terminate our ripALM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying
$\Delta^{k}_{\mathrm{res}}:=\max\left\{\Delta_{\rm kkt}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k}),\,\Delta_{\rm gap}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})\right\}<10^{-6}.$
Initial points. For the experiments in Section 5.1, we simply initialize all inexact pALMs at the origin to amplify the impact of inexactness. For the experiments in Section 5.2, we will employ a warm-start strategy for more efficiency. Indeed, our numerical experience (see, e.g., [21, 23, 51]) have suggested that a reasonably good initial point would benefit the practical performance of such a “pALM + Ssn” algorithmic framework. Therefore, for the experiments in Section 5.2, we first employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1) to generate an initial point for our ripALM; see Appendix C for more details. Specifically, we terminate this iBPGM as long as it produces a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ such that $\Delta^{k}_{\mathrm{res}}<10^{-3}$, or it reaches the maximal number of iterations 500. Note that the time consumed during the warm-starting phase is included in the total computational time for our ripALM. Additionally, the initial error variable $\bm{w}^{0}$ in our ripALM is always set to $\bm{0}$.

### v2

@@ 4.1 Quadratically regularized optimal transport
To apply the proposed ripALM, we first reformulate (4.1) as follows:
$\min_{X\in\mathbb{R}^{m\times n}}~~f_{\texttt{q}}(X):=\frac{\lambda}{2}\left\|{X}\right\|_{F}^{2}+\left\langle{C},{X}\right\rangle+\delta_{\mathbb{R}^{m\times n}_{+}}(X)\quad\mathrm{s.t.}\quad X\bm{1}_{n}=\bm{\alpha},~X^{\top}\bm{1}_{m}=\bm{\beta}.$ (4.2)
Clearly, it falls into the form of problem (1.1) (upon the vectorization of $X$). Further, one can show that the dual problem of (4.2) is given by (modulo a minus sign)
@@
and its associated augmented Lagrangian function is given by (using a similar deduction as Section 2)
@@
$X^{k+1}=\mathtt{prox}_{\sigma_{k}f_{\texttt{q}}}\left(X^{k}+\sigma_{k}\left(\bm{u}^{k+1}\bm{1}_{n}^{\top}+\bm{1}_{m}(\bm{v}^{k+1})^{\top}\right)\right),$ (4.5)
@@
To truly implement ripALM for solving problem (4.2), it is essential to efficiently solve the subproblem (4.4) to find $(\bm{u}^{k+1},\bm{v}^{k+1})$ satisfying the error criterion (2.5). In the following, we shall describe how to apply a semismooth Newton (Ssn) method to achieve this goal. For simplicity, we drop the index $k$ and explicitly rewrite the subproblem (4.4) as follows:
$\hskip-5.69054pt\min\limits_{\bm{u}\in\mathbb{R}^{m},\,\bm{v}\in\mathbb{R}^{n}}\left\{\begin{aligned} &\Psi(\bm{u},\bm{v}):=-\bm{\alpha}^{\top}\bm{u}-\bm{\beta}^{\top}\bm{v}+\frac{1}{2\sigma}\left\|\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right\|_{F}^{2}-\frac{1}{2\sigma}\|\bar{X}\|_{F}^{2}\\ &\quad-\mathtt{M}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)+\frac{\tau}{2\sigma}\left\|\bm{u}-\bar{\bm{u}}\right\|^{2}+\frac{\tau}{2\sigma}\left\|\bm{v}-\bar{\bm{v}}\right\|^{2}\end{aligned}\right\},$ (4.6)
@@
$\nabla\Psi(\bm{u},\bm{v})=\left[\begin{matrix}\left(\mathtt{prox}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)\right)\bm{1}_{n}-\bm{\alpha}+\tau\sigma^{-1}\left(\bm{u}-\bar{\bm{u}}\right)\vskip 2.84526pt\\ \left(\mathtt{prox}_{\sigma f_{\texttt{q}}}\left(\bar{X}+\sigma\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)\right)\right)^{\top}\bm{1}_{m}-\bm{\beta}+\tau\sigma^{-1}\left(\bm{v}-\bar{\bm{v}}\right)\end{matrix}\right].$
Note that the proximal mapping $\mathtt{prox}_{\sigma f_{\texttt{q}}}(\cdot)$ can be easily computed as follows:
$\mathtt{prox}_{\sigma f_{\texttt{q}}}\left(Z\right)=\frac{1}{1+\lambda\sigma}\Pi_{\mathbb{R}_{+}^{m\times n}}\left(Z-\sigma{C}\right),\quad\forall\,Z\in\mathbb{R}^{m\times n}.$
@@
In view of the nice property of $\nabla\Psi$, we are able to follow [22, 24, 25, 58] to apply a globally convergent and locally superlinearly convergent Ssn method to solve (4.7). Specifically, let $\mathcal{W}(\bm{u},\bm{v}):=\bar{X}+\sigma\left({\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}-C}\right)$ and define the multifunction $\widehat{\partial}(\nabla\Psi):\mathbb{R}^{m}\times\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{(m+n)\times(m+n)}$ as follows:
[…]
where $B:=\begin{bmatrix}\bm{1}_{n}^{\top}\otimes I_{m}\\ I_{n}\otimes\bm{1}_{m}^{\top}\end{bmatrix}\in\mathbb{R}^{(m+n)\times mn}$ with $\otimes$” denoting the Kronecker product, $\texttt{vec}(\Omega)$ denotes the vectorization of $\Omega$ with $\texttt{vec}(\Omega)]_{i+(j-1)m}=\Omega_{ij}$ for any $1\leq i\leq m$ and $1\leq j\leq n$, $\mathrm{Diag}(\bm{z})$ denotes the diagonal matrix whose $i$th diagonal element is given by $z_{i}$, and $\partial{\Pi_{\mathbb{R}_{+}^{m\times n}}}(Z)$ denotes the generalized Jacobian of the Lipschitz continuous mapping $\Pi_{\mathbb{R}_{+}^{m\times n}}$ at $Z$, which is defined by
@@
Then, using similar arguments as in the proof of [58, Proposition 4.1], one can show that $\nabla\Psi(\cdot)$ is strongly semi-smooth with respect to $\widehat{\partial}(\nabla\Psi)(\cdot)$, and thus the Ssn method is applicable. More importantly, for any $(\bm{u},\bm{v})\in\mathbb{R}^{m}\times\mathbb{R}^{n}$, all elements of $\widehat{\partial}(\nabla\Psi)(\bm{u},\bm{v})$ are positive definite. This ensures the direct applicability of the Ssn method without the need for a specific regularity condition, such as the primal constraint nondegeneracy condition [62], highlighting the advantage of incorporating a proximal term. The detailed description of the Ssn method for solving equation (4.7) is presented in Algorithm 2. We refer readers to [22, Theorem 3.6] for its detailed convergence results.
@@
4.2 Basis pursuit denoising
The basis pursuit denoising (BPDN) problem is a fundamental optimization problem in signal processing and compressed sensing, which aims to recover a sparse signal from noisy linear measurements [51]. Specifically, given a measurement $\bm{b}\in\mathbb{R}^{m}$, a dictionary $D\in\mathbb{R}^{m\times n}$, and an estimate of the noise level $\widehat{\kappa}>0$ in the data, the BPDN problem seeks a sparse solution $\bm{s}\in\mathbb{R}^{n}$ by solving the following constrained optimization problem:
$\min_{\bm{s}\in\mathbb{R}^{n}}~\|\bm{s}\|_{1}\quad\text{s.t.}\quad\|D\bm{s}-\bm{b}\|\leq\widehat{\kappa}$. (4.9)
Note that in the sparse recovery regime, one focuses on the case that $m\ll n$. By introducing an auxiliary variable $\bm{t}\in\mathbb{R}^{m}$, the above problem can be reformulated as
$\min_{\bm{s}\in\mathbb{R}^{n},\,\bm{t}\in\mathbb{R}^{m}}\;\|\bm{s}\|_{1}+\delta_{\mathcal{B}^{2}_{\widehat{\kappa}}}(\bm{t})\quad\text{s.t.}\quad D\bm{s}-\bm{t}=\bm{b},$ (4.10)
where $\mathcal{B}^{2}_{\widehat{\kappa}}\subset\mathbb{R}^{m}$ denotes the $\ell_{2}$-norm ball centered at $\bm{0}$ with radius $\widehat{\kappa}$. Then, we see that the reformulated problem (4.10) is a special case of the general problem (1.1) by letting $\bm{x}:=(\bm{s};\bm{t})\in\mathbb{R}^{n+m}$, $f(\bm{x}):=\|\bm{s}\|_{1}+\delta_{\mathcal{B}^{2}_{\widehat{\kappa}}}(\bm{t})$ and $A:=[D,-I]\in\mathbb{R}^{m\times(n+m)}$. Consequently, the proposed ripALM is directly applicable to solve the BPDN problem.
To apply ripALM, we derive the dual problem of (4.10) as follow (modulo a minus sign)
$\min_{\bm{y}\in\mathbb{R}^{m}}~~\delta_{\mathcal{B}_{1}^{\infty}}({D}^{\top}\bm{y})+\widehat{\kappa}\|\bm{y}\|-\bm{b}^{\top}\bm{y},$ (4.11)
where $\mathcal{B}_{1}^{\infty}\subset\mathbb{R}^{n}$ denotes the unit $\ell_{\infty}$-norm ball centered at $\bm{0}$. The associated augmented Lagrangian function is given by
$\mathcal{L}_{\sigma}(\bm{y},\bm{s},\bm{t})=$ $-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bm{s}+\sigma{D}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{s}\|^{2}-\mathtt{M}_{\sigma\|\cdot\|_{1}}\big(\bm{s}+\sigma{D}^{\top}\bm{y}\big)$
$+\frac{1}{2\sigma}\big\|\bm{t}-\sigma\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bm{t}\|^{2}-\mathtt{M}_{\sigma\delta_{\mathcal{B}_{\widehat{\kappa}}^{2}}}\big(\bm{t}-\sigma\bm{y}\big).$
Thus, given an arbitrary initial guess $(\bm{y}^{0},\bm{s}^{0},\bm{t}^{0})\in\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}^{m}$, the basic iterative scheme of ripALM for solving problem (4.11) reads as follows:
$\bm{y}^{k+1}\approx\arg\min_{\bm{y}\in\mathbb{R}^{m}}\left\{\mathcal{L}_{\sigma_{k}}(\bm{y},\bm{s}^{k},\bm{t}^{k})+\frac{\tau_{k}}{2\sigma_{k}}\|\bm{y}-\bm{y}^{k}\|^{2}\right\},$ (4.12)
$\bm{s}^{k+1}=\mathtt{prox}_{\sigma_{k}\|\cdot\|_{1}}\big(\bm{s}^{k}+\sigma_{k}D^{\top}\bm{y}^{k+1}\big),$ (4.13)
$\bm{t}^{k+1}=\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}\big(\bm{t}^{k}-\sigma_{k}\bm{y}^{k+1}\big),$ (4.14)
where $\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}$ denotes the projection operator onto $\mathcal{B}_{\widehat{\kappa}}^{2}$.
To truly implement ripALM for solving problem (4.11), it is essential to efficiently solve the subproblem (4.12) to find $\bm{y}^{k+1}$ satisfying the relative error criterion (2.5). In the following, we describe how to apply the Ssn method to achieve this goal. For simplicity, we drop the index $k$ and explicitly rewrite the subproblem (4.12) as follows:
$\min_{\bm{y}\in\mathbb{R}^{m}}~~\left\{\begin{aligned} \Phi(\bm{y}):=&-\bm{b}^{\top}\bm{y}+\frac{1}{2\sigma}\big\|\bar{\bm{s}}+\sigma{D}^{\top}\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bar{\bm{s}}\|^{2}-\mathtt{M}_{\sigma\|\cdot\|_{1}}\big(\bar{\bm{s}}+\sigma{D}^{\top}\bm{y}\big)\\ &+\frac{1}{2\sigma}\big\|\bar{\bm{t}}-\sigma\bm{y}\big\|^{2}-\frac{1}{2\sigma}\|\bar{\bm{t}}\|^{2}-\mathtt{M}_{\sigma\delta_{\mathcal{B}^{2}_{\widehat{\kappa}}}}\big(\bar{\bm{t}}-\sigma\bm{y}\big)+\frac{\tau}{2\sigma}\|\bm{y}-\bar{\bm{y}}\|^{2}\end{aligned}\right\}.$
where $\bar{\bm{s}}$, $\bar{\bm{t}}$ and $\bar{\bm{y}}$ are given vectors. From the property of the Moreau envelope, we see that $\Psi$ is strongly convex and continuously differentiable with the gradient
$\nabla\Phi(\bm{y})={D}\mathtt{prox}_{\sigma\|\cdot\|_{1}}\big(\bar{\bm{s}}+\sigma D^{\top}\bm{y}\big)-\Pi_{{\mathcal{B}_{\widehat{\kappa}}^{2}}}\big(\bar{\bm{t}}-\sigma\bm{y}\big)-\bm{b}+\frac{\tau}{\sigma}\big(\bm{y}-\bar{\bm{y}}\big).$ (4.15)
$\nabla\Phi(\bm{y})=0.$ (4.16)
Next, we define the multifunction $\widehat{\partial}(\nabla\Phi):\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ as follows:
[…]
where $\partial\mathtt{prox}_{\sigma\|\cdot\|_{1}}:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n\times n}$ is defined by
[…]
and $\partial\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}:\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m\times m}$ is defined by
[…]
Using the relation $\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}(\bm{x})=\bm{x}-\mathtt{prox}_{\widehat{\kappa}\|\cdot\|_{2}}(\bm{x})$, Lemma 2.1 in [60], and similar arguments to those in [58, Proposition 4.1], one can show that $\nabla\Phi$ is strongly semismooth with respect to the multifunction $\widehat{\partial}(\nabla\Phi)$. This property enables the application of the Ssn method to solve (4.16). Since the algorithmic framework is identical to Algorithm 2, we omit the detailed description for simplicity. Moreover, note from (4.17) that, for any $\bm{y}\in\mathbb{R}^{m}$, all matrices in $\widehat{\partial}(\nabla\Phi)(\bm{y})$ are symmetric positive definite. This ensures the local superlinear (even quadratic) convergence rate of the Ssn method, without requiring additional regularity conditions. These favorable theoretical properties form the basis for the encouraging numerical performance, as observed from our numerical results in the next section.
Finally, we would like to point out that the main computational cost of the Ssn method for solving (4.16) lies in solving a sequence of linear systems of the form
$H\bm{d}=-\nabla\Phi(\bm{y}),$ (4.18)
to obtain the Newton direction $\bm{d}$, where $H\in\widehat{\partial}(\nabla\Phi)(\bm{y})$ with $\widehat{\partial}(\nabla\Phi)$ defined in (4.17). Inspired by the techniques developed in [22, Section 3.3], we can exploit the special structure of $H$ to enhance the computational efficiency of solving (4.18). However, the presence of the additional block $V$ in $H$ introduces nontrivial implementation difficulties. Detailed discussions on efficiently solving (4.18) are provided in Appendix D.
@@ 5 Numerical experiments
In this section, we conduct numerical experiments to evaluate the performance and validate the efficiency of our ripALM in Algorithm 1 for solving the QROT problem (4.1) and the BPDN problem (4.9). Specifically, we conduct experiments from the following three aspects:
  • In Section 5.1, we compare our ripALM with Snipal [24] and ciPALM [58], which represent the latest absolute-type and relative-type inexact pALMs, respectively, for solving the QROT problem on image data. These comparisons aim to illustrate how different error criteria with different tolerance parameters influence the practical numerical performance of the inexact pALM.
  • In Section 5.2, we compare our ripALM with Gurobi, Eckstein and Silva’s method [14] (denoted by ES-iALM) and a dual alternating direction method of multipliers (dADMM) presented in Appendix B, for solving large-scale QROT problems. Gurobi is one of the state-of-the-art commercial solvers, while ES-iALM is a relative-type inexact version of the vanilla ALM. Moreover, dADMM is a widely used first-order method for large-scale constrained convex optimization; see, for example, [6, 17]. In addition, it is worth noting that many other first-order methods can also be applied to the QROT problem, such as the Chambolle–Pock algorithm [8], the Condat–Vũ algorithm [9, 52], and the linearized ALM [53, 55]. However, our preliminary tests show that these methods often perform comparably to, or worse than, dADMM. For this reason, and for the sake of simplicity, we omit them from our comparisons and only select the dADMM in Appendix B as a representative first-order method.
  • In Section 5.3, we compare our ripALM with SPGL1 [51], Mosek, and the dADMM presented in Appendix E, for solving the BPDN problem. SPGL1 is a state-of-the-art solver specifically designed for BPDN, while Mosek is a highly efficient commercial solver that can be applied to solve the second-order cone programming reformulation of the BPDN problem; see Appendix F for more details.
@@
Termination conditions. We first describe the termination conditions for QROT. Let $\mathcal{Z}(\bm{u},\bm{v},X):=C+\lambda X-\bm{u}\bm{1}_{n}^{\top}-\bm{1}_{m}\bm{v}^{\top}$. The Karush-Kuhn-Tucker (KKT) system for problem (4.1) and its dual (4.3) is given by
@@
$\Delta_{\rm kkt}^{\rm qrot}(\bm{u},\bm{v},X):=\max\left\{\Delta_{p}^{\rm qrot}(X),\,\Delta_{d}^{\rm qrot}(\bm{u},\bm{v},X),\,\Delta_{c}^{\rm qrot}(\bm{u},\bm{v},X)\right\},$
@@
$\;\Delta_{p}^{\rm qrot}(X):=\max\left\{\frac{\|X\bm{1}_{n}-\bm{\alpha}\|}{1+\|\bm{\alpha}\|},\,\frac{\|X^{\top}\bm{1}_{m}-\bm{\beta}\|}{1+\|\bm{\beta}\|},\,\frac{\|\min\{X,\,0\}\|_{F}}{1+\|X\|_{F}}\right\},$
$\;\Delta_{d}^{\rm qrot}(\bm{u},\bm{v},X):=\frac{\|\min\{\mathcal{Z}(\bm{u},\,\bm{v},\,X),\,0\}\|_{F}}{1+\|C\|_{F}},\quad\;\Delta_{c}^{\rm qrot}(\bm{u},\bm{v},X):=\frac{\left|\langle X,\,\mathcal{Z}(\bm{u},\,\bm{v},\,X)\rangle\right|}{1+\|C\|_{F}}.$
@@
$\Delta_{\rm gap}^{\rm qrot}(\bm{u},\,\bm{v},\,X):=\cfrac{\left|\mathtt{pobj}(X)-\mathtt{dobj}(\bm{u},\bm{v})\right|}{1+\left|\mathtt{pobj}(X)\right|+\left|\mathtt{dobj}(\bm{u},\bm{v})\right|},$
where $\mathtt{pobj}(X):=f_{\texttt{q}}(X)$ and $\mathtt{dobj}(\bm{u},\bm{v}):=-f_{\texttt{q}}^{*}\left(\bm{u}\bm{1}_{n}^{\top}+\bm{1}_{m}\bm{v}^{\top}\right)+\bm{\alpha}^{\top}\bm{u}+\bm{\beta}^{\top}\bm{v}$. Using these relative residuals, we will terminate ripALM when it returns a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ satisfying
$\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k}):=\max\left\{\Delta_{\rm kkt}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k}),\,\Delta_{\rm gap}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})\right\}<10^{-6}.$
Similarly, we set up the termination conditions for BPDN based on the associated KKT system. Specifically, given an output solution $(\bm{s},\bm{t},\bm{y})$, we define
$\Delta_{p}^{\rm bpdn}(\bm{s},\bm{t})$ $:=\cfrac{\left\|D\bm{s}-\bm{b}-\bm{t}\right\|}{1+\|\bm{b}\|},$
$\Delta_{d}^{\rm bpdn}(\bm{s},\bm{t},\bm{y})$ $:=\cfrac{\sqrt{\big\|\bm{s}-\mathtt{prox}_{\|\cdot\|_{1}}\left(\bm{s}+{D}^{\top}\bm{y}\right)\big\|^{2}+\big\|\bm{t}-\Pi_{\mathcal{B}_{\widehat{\kappa}}^{2}}\left(\bm{t}-\bm{y}\right)\big\|^{2}}}{1+\|D\|_{F}},$
${\Delta}_{\rm gap}^{\rm bpdn}(\bm{s},\bm{t},\bm{y})$ $:=\cfrac{\left|\mathtt{pobj}(\bm{s},\bm{t})-\mathtt{dobj}(\bm{y})\right|}{1+\left|\mathtt{pobj}(\bm{s},\bm{t})\right|+\left|\mathtt{dobj}(\bm{y})\right|},$
where $\mathtt{pobj}(\bm{s},\bm{t}):=\|\bm{s}\|_{1}$ and $\mathtt{dobj}(\bm{y}):=-\widehat{\kappa}\|\bm{y}\|+\bm{b}^{\top}\bm{y}$. Then, we terminate ripALM when it returns $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ such that
$\Delta_{\rm res}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k}):=\max\left\{\Delta_{p}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k}),\,\Delta_{d}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k}),\,\Delta_{\rm gap}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})\right\}<10^{-6}.$
Initial points. For the experiments in Section 5.1, we simply initialize all inexact pALMs at the origin to amplify the impact of inexactness. For the experiments in Section 5.2 and Section 5.3, we will employ a warm-start strategy for more efficiency. Indeed, our numerical experience (see, e.g., [24, 26, 58]) have suggested that a reasonably good initial point would benefit the practical performance of such a “pALM + Ssn” algorithmic framework. Therefore, for the experiments in Section 5.2, we first employ an inexact Bregman proximal gradient method (iBPGM) with Sinkhorn’s algorithm as a subsolver for solving the QROT problem (4.1) to generate an initial point for our ripALM; see Appendix C for more details. Specifically, we terminate this iBPGM as long as it produces a point $(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})$ such that $\Delta_{\mathrm{res}}^{\rm qrot}(\bm{u}^{k},\,\bm{v}^{k},\,X^{k})<10^{-3}$, or its iteration number reaches 500. Moreover, we also apply this warm-start strategy to ES-iALM in Section 5.2 to ensure a fair comparison. For the experiments in Section 5.3, we use dADMM in Appendix E to generate an initial point. Specifically, we terminate dADMM as long as it produces a point $(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})$ such that $\Delta_{\mathrm{res}}^{\rm bpdn}(\bm{s}^{k},\bm{t}^{k},\bm{y}^{k})<10^{-3}$, or its iteration number reaches 100. Note that the time consumed during the above warm-starting phase is included in the total computational time for both ripALM and ES-iALM. In addition, the initial error variable $\bm{w}^{0}$ in ripALM and ES-iALM is always set to $\bm{0}$.